# Controlled bit-flip of period-doubling and discrete time crystalline states in open systems

### v1

@@ III.1 Ideal limit without noise
We present in Figs. 2(c)-2(f) the zero-temperature bit-flip diagram of the DPP for increasing $\gamma$. The bit-flip diagram identifies the parameter regimes where we can successfully flip a bit encoded in the PD state. Here, $\delta A=A-A_{r}$ corresponds to the detuning of the driving amplitude from $A_{r}$. We also show in Figs. 2(g)-2(j) the same set of bit-flip diagrams for the ODM for increasing $\kappa$ and $\lambda_{0}=0.9\lambda_{c}$ in the thermodynamic limit.
To determine whether we can switch a PD or DTC state at a given $\delta A$ and $T_{\delta}$, we first initialize the systems in their stable fixed points at $t_{i}<0$. For the DPP, we consider the initial state
@@
with $\epsilon=10^{-6}$. We then apply a periodic drive to push the DPP (ODM) into a PD (DTC) state for 100 driving cycles. At $t=0$, we finally apply the defect protocol in Eq. (14) until $t=T_{\delta}$, as exemplified in Fig. 1(b). We then allow the system to relax back to a new PD (DTC) state before obtaining the difference between the steady-state values of the absolute-time phase before and after the defect, $\triangle\varphi$. If $\triangle\varphi=\pi$, we have a successful bit-flip between the PD and DTC states, while we have no bit-flip when $\triangle\varphi=0$. Note that in evaluating $\triangle\varphi$, we enforced the periodic boundary of the absolute-time phase such that $\varphi\in[-\pi,\pi)$.
While the bit-flip diagrams of the DPP and the ODM in Fig. 2 appears to be quantitatively distinct from each other, there is a clear qualitative trend for varying dissipation strengths. For weak dissipation, for instance, the bit-flip diagram exhibits complex, fractal-like structures. As we further increase the dissipation strengths, the fractality in the diagrams eventually vanishes until they become uniform for sufficiently strong dissipation over a wide range of $\delta A$ and $T_{\delta}$. This behavior suggests that in the zero-temperature limit and the thermodynamic limit for the DPP and ODM, respectively, bit-flip operations become more robust when the dissipation is strong, in that perturbations in $A$ and $T_{\delta}$ will not significantly affect the success rate of the bit-flip between the states. This is in contrast to weak dissipation, where there is more uncertainty in the bit-flip success due to the fractal structures in the bit-flip diagram.
where $\vec{s}_{\mathrm{O},\mathrm{P}}=\begin{pmatrix}S^{x}_{\mathrm{O},\mathrm{P}}&S^{y}_{\mathrm{O},\mathrm{P}}&S^{z}_{\mathrm{O,P}}\end{pmatrix}^{T}/N$, and by the conservation of spin angular momentum for the ODM [70], $||\vec{s}_{O}||^{2}=||\vec{s}_{P}||^{2}=1/4$. The decorrelator determines the deviation between two identical systems, labeled as systems $\mathrm{O}$ and $\mathrm{P}$, initialized in a nearly identical state [39]. In particular, for regular dynamics, we expect that $\vec{s}_{\mathrm{O}}$ and $\vec{s}_{\mathrm{P}}$ will be aligned at all times and thus $d^{2}(t)=0$. Meanwhile, we get the maximum $d^{2}=1$ when $\vec{s}_{\mathrm{O}}$ and $\vec{s}_{\mathrm{P}}$ are anti-aligned with one another.
Given that strong dissipation allows for a more robust bit-flip between PD and DTC states, we will determine in the next subsection whether the bit-flip operation remains robust even in the presence of thermal and quantum fluctuations for the DPP and ODM, respectively.
@@ III.2 Effects of thermal and quantum noise
To explore the robustness of the bit-flip operation of PD states against thermal fluctuations, we numerically integrate the Langevin equation in Eq. (1) using a predictor-corrector method, with a time step of $\triangle t=0.01\Omega$. The initial state is the same as in the zero-temperature limit given by Eq. (16). We find that high temperatures destroy the coherence of the parametric oscillation in the long-time limit (see Appendix B.2 for details), and as such, we only consider a range of temperature from $\tilde{T}=10^{-6}$ to $\tilde{T}=2\times 10^{-4}$.
For the DTCs in the ODM at zero temperature, we explore the effects of quantum noise on the robustness of bit-flip operations by approximating the solutions of Eq. (7) in two ways. We first consider the truncated Wigner approximation (TWA), in which, similar to the mean-field equations, we treat the operators as complex numbers. The quantum fluctuations, in this case, come from two sources: the fluctuations in the initial states of the cavity mode and the temporal fluctuations due to the dissipation, $\xi(t)$ [74, 69]. In this approach, we initialize the cavity mode in a vacuum state, which is numerically represented as complex Gaussian variable $a_{0}=\frac{1}{2}\left(\zeta_{\mathrm{R}}+i\zeta_{\mathrm{I}}\right)$, where $\zeta_{\mathrm{R},\mathrm{I}}$ are random numbers sampled from a Gaussian distribution satisfying the following conditions,
@@
for $i,j=\mathrm{R},\mathrm{I}$ [75]. We also consider a second approach using the discrete truncated Wigner approximation (DTWA), in which the quantum noise is included in the individual spins rather than treating them as a single collective spin. In this approach, each initial spin state has a nonzero $z$-component that is oriented at the $-z$ direction. Each of their $x$ and $y$ components however are randomly sampled according to their respective Wigner distribution [76, 77]. Using this approach, we can further assess the effects of finite size on the robustness of DTC switching in our system. The corresponding equations of motion for both the TWA and the DTWA are shown in Appendix A, and we integrate them using the same predictor-corrector method used for the DPP with a time step of $\Delta t=0.01\omega$
Figure 4: (a) Switching probability of the resonantly-driven DPP as a function of $T_{\delta}$ for $\gamma=0.1\Omega$, $\delta A=0.4$ and different values of $\tilde{T}$. (b)-(c) Switching probability of the resonantly-driven ODM for (b) different $N$ and (c) different set of stochastic noise for $N=5\times 10^{3}$. For (b), the dynamics are obtained using the TWA. The remaining parameters are $\lambda_{0}=0.9\lambda_{c}$, $\kappa=1.0\omega$ and $\delta A=0.1$. For both systems, we considered $M=1000$ trajectories.
We present in Fig. 4(a) the success probability of a bit flip for $M=1000$ trajectories, $P_{s}$, for the PD states as a function of $T_{\delta}$ for $\tilde{T}\neq 0$. We can observe that for very small $\tilde{T}$, $P_{s}$ has a discontinuous transition from $P_{s}=0$ to $P_{s}=1$ at $T_{\delta,c}\approx 10T_{d}$, akin to those observed in Ref. [8]. As we increase the temperature, however, the thermal fluctuations soften the discontinuity into a crossover. This allows the system to have a nonzero probability to switch between PD states at lower $T_{\delta}$. We can see a similar effect of quantum noise on the DTCs for finite $N$, as depicted in Fig. 4(b). As expected, the transition between unsuccessful and successful bit-flip operation becomes sharper as $N$ increases and approaches the shape of that in the thermodynamic limit, $N\rightarrow\infty$. This can be understood from the inverse proportionality between the temporal noise strength and the particle number as highlighted in Eq. (23).
The quantum noise on the ODM originates from two sources: (i) initial quantum fluctuations and (ii) temporal fluctuations due to the dissipation. As such, it is natural to ask how these two types of fluctuations contribute to the crossover-like behavior of $P_{s}$. To this end, we investigate the $P_{s}$ of the ODM when there is only initial noise in the cavity and spin degrees of freedom. We present this in Fig. 4(c) together with the $P_{s}$ obtained using the TWA and the DTWA. Notably, when only initial noise is present on the cavity and individual spins, we observe the discontinuous transition seen in the large $N$ limit, which is in contrast to the crossover-like feature of $P_{s}$ in the TWA and DTWA. We attribute this to dissipation removing any effects of the initial noise on the dynamics in the absence of temporal fluctuations, leading to an effectively mean-field dynamics on a single-trajectory level, and thus the mean-field characteristics of $P_{s}$.
In Fig. 4(b), unlike the crossover in Fig. 4(a), the $P_{s}$ of the ODM for $N=5\times 10^{3}$ and $N=10^{4}$ exhibits a plateau within the interval $T_{\delta}/T_{d}\in\left[1.0,1.5\right]$ before approaching $P_{s}=1$ for large $T_{\delta}$. This effect is more apparent in Fig. 4(c), wherein the $P_{s}$ approaches a plateau at $P_{s}\approx 0.75$. To better understand the appearance of such plateaus, we further characterize the switching dynamics of the PD and DTC states using the half winding number $w$, introduced in Ref. [8] to characterize the dynamics of a PD state during the defect. It is defined as [8, 63],
[…]
@@
and it quantifies the number of times the $O_{R}(t)$ performs a half-rotation around the origin of the complex plane. To illustrate this for a DTC during a defect protocol, we present in Figs. 5(a)-5(c) exemplary dynamics of $X=\mathrm{Re}[O_{R}]$ and $Y=\mathrm{Im}[O_{R}]$. Here, we rescale the axes by the maximum value of $|R(t)|$, $R_{m}$. Similar to the results in Ref. [8], and Figs. 1(d), 1(f), and 1(h), during the defect protocol, the trajectory of the $O_{R}$ approaches the origin. The half-winding number depends on whether $O_{R}$ rotates around the origin or not. If $O_{R}$ avoids the origin, then $w=0$, and thus the PD or DTC state does not switch. On the other hand, if $O_{R}$ does a single half counter-clockwise rotation around the origin, then $w=+1$, meaning the PD or DTC state switches to its symmetry broken partner. The same thing occurs when $w=-1$, except that $O_{R}$ now rotates in a clockwise manner. Note that we only get a bit-flip when $w=2n+1$, with $n\in\mathbb{Z}$, since an even-valued $w$ implies a complete rotation around the origin and that the system returns to its original state. Note that in our convention for obtaining $\varphi(t)$ using Eq. (6), the half-winding number of the DPP and the ODM in the limit of $T_{\delta}\gg T_{d}$ is $w=-1$.
@@
Figure 5: (a)-(c) Exemplary dynamics of the complex response amplitude of $S^{x}/N$ for (a) $w=1$, (b) $w=0$, and (c) $w=-1$. The corresponding defect durations are (a) $T_{\delta}=0.95T_{d}$, (b) $T_{\delta}=0.5T_{d}$, and (c) $T_{\delta}=1.5$, respectively. (d)-(g) Half-winding number distribution of the resonantly-driven DPP for (d) $\tilde{T}=2\times 10^{-4}$, (e) $\tilde{T}=10^{-4}$, (f) $\tilde{T}=10^{-5}$, and (g) $\tilde{T}=10^{-6}$. (h)-(k) Half-winding number distribution of the ODM at resonance for (h) $N=5\times 10^{3}$, (i) $N=10^{4}$, (j) $N=10^{5}$, (k) $N=10^{6}$. The remaining parameters for the thermal DPP are $\gamma=0.1\Omega$, and $\delta A=0.4$, while the remaining parameters for the ODM are $\lambda_{0}=0.9\lambda_{c}$, $\kappa=1.0\omega$, and $\delta A=0.1$.
@@
We now show in Figs. 5(d)-5(g) the probability distribution $P_{w}$ of $w$ as a function of $T_{\delta}$ and $\tilde{T}$ for the thermal DPP. Note that we have used the same parameters considered in Fig. 4(a). We can see that the $P_{w}$ is localized in a small subset of $w$ when $\tilde{T}\ll 1$. Specifically, for $T_{\delta}<T_{\delta,c}$, $w=0$, and thus no bit-flip occurs. Whereas for $T_{\delta}>T_{\delta,c}$, $w=-1$, which means that the PD state switches as a result of the defect protocol. We can also observe that $w$ is discontinuous exactly at $T_{\delta}=T_{\delta,c}$, consistent with the behavior observed in Ref. [8]. However, as we increase $\tilde{T}$, the discontinuity at $T_{\delta,c}$ becomes blurry as $P_{w}$ spreads out such that the DPP can either have $w=0$ or $w=-1$ near $T_{\delta,c}$, i.e. it is now uncertain if the bit-flip operation will work or not. This behavior is consistent with the crossover observed in Fig. 4(a). Note that in the case of $\tilde{T}=2\times 10^{-4}$, as shown in Fig. 5(d), the faint $w=+1$ branch in the $P_{w}$ for $T_{\delta}<T_{\delta,c}$ is due to thermally-activated switching between the PD states, in which due to the finite coherence time of the oscillations, the thermal noise can become strong enough to flip a PD state [78].
@@
The situation is different in the case of the $P_{w}$ of the ODM, as shown in Figs. 5(h)-5(k). While we still observe qualitatively similar behavior of the distribution as in the DPP, such as the localization of $P_{w}$ into one value of $w$ as $N$ increases, the $P_{w}$ of the ODM also reveals a $w=+1$ branch from $T_{\delta}\approx 0.9T_{d}$ to $T_{\delta}=1.0T_{d}$. In Fig. 5(k), we show that when $N$ is sufficiently large, the $P_{w}$ in these regions of $T_{\delta}$ becomes more concentrated into $w=+1$, which explains the discontinuous transition at $T_{\delta}\approx 0.9T_{d}$ shown in Fig. 4(b). As we decrease $N$, however, the system obtains nonzero probability to either have a half-winding number of $w=\pm 1$ or $w=0$, thus resulting in the plateau observed in Figs. 4(b) and 4(c). Note that while we do not observe any appearance of plateaus in the thermal DPP for the parameters considered here, we still expect this behavior to appear when the system has a transition between $w=-1$ and $w=+1$, and the range of $T_{\delta}$ in which one of the branches exists is small enough for the system to have nonzero probability to get either $w=0$ or $w=\pm 1$.
@@
Our results imply two things about the robustness of the bit-flip operation beyond the weak-dissipation and noiseless limit. First, in the noiseless limit, strong dissipation enhances the robustness of the operation against perturbations in the driving and defect parameters. This leads to a wider range of $A$ and $T_{\delta}$ wherein controlled switching between PD and DTC states can be observed. Second, in the presence of fluctuation for fixed dissipation, not only does the PD and DTC switching remain robust for large $T_{\delta}$, but the noise also enhances the switching probability for $T_{\delta}<T_{\delta,c}$. Thus, we demonstrate that the phase ramp as a defect is a robust method for performing bit-flips between the PD states of thermal DPP and the DTC states of finite-sized ODM for $\lambda_{0}<\lambda_{c}$.
@@
Given that the DTC states of the ODM for $\lambda_{0}<\lambda_{c}$ can be mapped onto a PD state of a coupled oscillator model [62], we now explore in the next subsection whether we can still switch DTC states using the defect protocol in Eq. (14) when $\lambda_{0}>\lambda_{c}$, in which the mapping of the ODM onto the DPP breaks down.

### v2

@@ III.1 Ideal limit without noise
Before we proceed, let us first establish the ideal noiseless limit for the PO and the ODM. From Eq. (3), we can observe that the PO do not experience any stochastic noise when $\tilde{T}=0$ or when $\gamma=0$. As such, throughout this subsection, we will consider the zero-temperature limit and allow the dissipation $\gamma$ to be a control parameter. As for the ODM, as discussed in Appendix A, we reach the noiseless limit in the thermodynamic limit where the quantum fluctuations are suppressed. In this case, the dynamics of the system is described by the mean-field equations given in Appendix A.
We present in Figs. 2(c)–2(f) the zero-temperature bit-flip diagram of the PO for increasing $\gamma$. The bit-flip diagram identifies the parameter regimes where we can successfully flip a bit encoded in the period-doubled state. Here, $\delta A=A-A_{r}$ corresponds to the detuning of the driving amplitude from $A_{r}$. We also show in Figs. 2(g)-2(j) the same set of bit-flip diagrams for the ODM for increasing $\kappa$ and $\lambda_{0}=0.9\lambda_{c}$ in the thermodynamic limit.
To determine whether we can switch a period-doubled or DTC state at a given $\delta A$ and $T_{\delta}$, we first initialize the systems in their stable fixed points at $t_{i}<0$. For the PO, we consider the initial state
@@
where $a\equiv\langle\hat{a}\rangle$ and $S^{x,y,z}=\langle\hat{S}^{x,y,z}\rangle$ with $\epsilon=10^{-6}$. We then apply a periodic drive to push the PO (ODM) into a period-doubled (DTC) state for 100 driving cycles. At $t=0$, we finally apply the defect protocol in Eq. (12) until $t=T_{\delta}$, as exemplified in Fig. 1(b). We then allow the system to relax back to a new period-doubled (DTC) state before obtaining the difference between the steady-state values of the absolute-time phase before and after the defect, $\triangle\varphi$. If $\triangle\varphi=\pi$, we have a successful bit-flip, while we have no bit-flip when $\triangle\varphi=0$. Note that in evaluating $\triangle\varphi$, we enforced the periodic boundary of the absolute-time phase such that $\varphi\in[-\pi,\pi)$.
While the bit-flip diagrams of the PO and the ODM in Fig. 2 appear to be quantitatively distinct from each other, there is a clear qualitative trend for varying dissipation strengths. For weak dissipation, for instance, the bit-flip diagram exhibits complex, fractal-like structures. These fractal structures, as discussed in Appendix C, emerge due to the transient irregular dynamics induced by the defect protocol. These transient irregular dynamics becomes suppressed, however, by strong dissipation, resulting to the loss of fractal structures at large dissipation strengths. This behavior suggests that in the zero-temperature limit and the thermodynamic limit for the PO and ODM, respectively, bit-flip operations become more robust when the dissipation is strong, in that perturbations in $A$ and $T_{\delta}$ will not significantly affect the success rate of the bit-flip between the states. This is in contrast to weak dissipation, where there is more uncertainty in the bit-flip success due to the fractal structures in the bit-flip diagram.
Given that strong dissipation allows for a more robust bit-flip between period-doubled and DTC states, we will determine in the next subsection whether the bit-flip operation remains robust even in the presence of thermal and quantum fluctuations for the PO and ODM, respectively.
@@ III.2 Effects of thermal and quantum noise
We explore the robustness of the bit-flip operation of period-doubled states against thermal noise by numerically integrating the Langevin equation in Eq. (1) using a predictor-corrector method, with a time step of $\Omega\triangle t=0.01$. The initial state is the same as in the zero-temperature limit given by Eq. (14). We find that high temperatures destroy the coherence of the parametric oscillation in the long-time limit (see Appendix B.2 for details), and as such, we only consider a range of temperature from $\tilde{T}=10^{-6}$ to $\tilde{T}=2\times 10^{-4}$.
As for the DTCs in the ODM, a way to account the quantum fluctuations is to explicitly solve the Lindbland master equation in Eq. (7) to obtain the density matrix $\hat{\rho}(t)$ of the system. While this gives the exact quantum dynamics of the ODM, this method is computationally expensive for large $N$ since the Hilbert space of the system grows exponentially with $N$. To circumvent this problem, we employ an approximation called the truncated Wigner approximation (TWA) and its counterpart for spin systems, the discrete truncated Wigner approximation (DTWA) to solve the dynamics of the system. The TWA and DTWA work by evolving an ensemble of trajectories with initial values sampled from the Wigner distributions of the spin and cavity modes, and extracting the statistics of an observable from the ensemble [63, 64, 65]. In particular, for both the TWA and DTWA, we treat the cavity mode $\hat{a}$ as a complex variable $a(t)$ with a vacuum state as its initial state. This can be represented as a complex Gaussian variable $a(t_{i})=\left(\zeta_{\mathrm{R}}+i\zeta_{\mathrm{I}}\right)/2$ satisfying the conditions $\langle\zeta_{i}\rangle=0$ and $\langle\zeta_{i}\zeta_{j}\rangle=\delta_{i,j}$ for $i,j\in\left\{\mathrm{R},\mathrm{I}\right\}$ [66]. As for the spin component, in TWA, we treat the collective spin operators $\hat{S}^{x,y,z}$ as real variables with initial states given by Eq. (15). Whereas for the DTWA, we treat each two-level system $\hat{\sigma}_{i}^{x,y,z}$ as separate real variables, with $S^{x,y,z}=\sum_{i=1}^{N}\sigma_{i}^{z,y,z}/2$. In this case, we initialize the $z$-component of the individual spins at $\sigma^{z}_{i}=-1$, while we sample the $x$ and $y$ components of the spins from the set $(\sigma_{i}^{x},\sigma_{i}^{y})\in\left\{(1,1),(1,-1),(-1,1),(-1,-1)\right\}$ with equal probability. Thus, the DTWA captures the quantum noise contributed by each two-level system.
Both the TWA and DTWA account for the temporal noise neglected in the mean-field equations by including a stochastic term in the system’s equations of motion. In particular, given that the ODM is connected to a Markovian bath, the dynamics of the cavity and spin components can be described instead by the Heisenberg-Langevin equation [67],
$\partial_{t}\hat{a}=i\left[\hat{H}/\hbar,\hat{a}\right]-\kappa\hat{a}+\hat{\xi}(t),$ (16a)
$\partial_{t}\hat{S}^{x,y,z}=i\left[\hat{H}/\hbar,\hat{S}^{x,y,z}\right],$ (16b)
$\partial_{t}\hat{\sigma}_{i}^{x,y,z}=i\left[\hat{H}/\hbar,\hat{\sigma}_{i}^{x,y,z}\right],$ (16c)
where $\hat{\xi}(t)$ is a Gaussian noise operator satisfying the conditions,
$\hat{\xi}(t)\right>=0,\quad\left<\hat{\xi}(t)\hat{\xi}(t^{\prime})\right>=\kappa\delta\left(t-t^{\prime}\right).$ (17)
The noise term $\hat{\xi}$ can be viewed as fluctuations due to the photons entering the system from the Markovian bath as shown in Fig. 2(b). Mathematically, $\hat{\xi}(t)$ preserves the commutation relation of the cavity mode after evaluating the commutations in Eq. (16) and replacing the operators $\hat{a}$, $\hat{S}^{x,y,z}$, and $\sigma^{x,y,z}$, with the complex variables $a$, $S^{x,y,z}$, and $\sigma^{x,y,z}$, respectively. Following this procedure results to the stochastic equations of motions shown in Appendix A. With these equations of motion, we can calculate the system’s dynamics for an ensemble of initial states and obtain the statistics of the quantities we are interested in. For both the TWA and DTWA, we solve the equations of motion using the predictor-corrector method with a time step of $\omega\triangle t=0.01$
Figure 3: (a) Switching probability of the resonantly driven PO as a function of $T_{\delta}$ for $\gamma=0.1\Omega$, $\delta A=0.4$, and different values of $\tilde{T}$. (b), (c) Switching probability of the resonantly driven ODM for (b) different $N$ and (c) different set of stochastic noise for $N=5\times 10^{3}$. For (b), the dynamics are obtained using the TWA. The remaining parameters are $\lambda_{0}=0.9\lambda_{c}$, $\kappa=1.0\omega$, and $\delta A=0.1$. For both systems, we considered $1000$ trajectories.
We present in Fig. 3(a) the success probability of a bit flip between period-doubled states, $P_{s}$, as a function of $T_{\delta}$ for $\tilde{T}\neq 0$ and $1000$ trajectories. We can observe that for very small $\tilde{T}$, $P_{s}$ has a discontinuous transition from $P_{s}=0$ to $1$ at $T_{\delta,c}\approx 10T_{d}$, akin to those observed in Ref. [8]. As we increase the temperature, however, the thermal fluctuations soften the discontinuity into a crossover. This allows the system to have a nonzero probability to switch between period-doubled states at lower $T_{\delta}$. We can see a similar effect of quantum noise on the DTCs for finite $N$, as depicted in Fig. 3(b). As expected, the transition between unsuccessful and successful bit-flips becomes sharper as $N$ increases and approaches the shape of that in the thermodynamic limit, $N\rightarrow\infty$. This can be understood from the inverse proportionality between the temporal noise strength and the particle number as highlighted in Eq. (21).
The quantum noise on the ODM originates from two sources: (i) initial quantum fluctuations and (ii) temporal fluctuations due to the dissipation. As such, it is natural to ask how these two types of fluctuations contribute to the crossover-like behavior of $P_{s}$. To this end, we investigate the $P_{s}$ of the ODM when there is only initial noise in the cavity and spin degrees of freedom. We present this in Fig. 3(c) together with the $P_{s}$ obtained using the TWA and the DTWA. Notably, when only initial noise is present on the cavity and individual spins, we observe the discontinuous transition seen in the large-$N$ limit, which is in contrast to the crossover-like feature of $P_{s}$ in the TWA and DTWA. We attribute this to dissipation removing any effects of the initial noise on the dynamics in the absence of temporal fluctuations, leading to an effectively mean-field dynamics on a single-trajectory level, and thus the mean-field characteristics of $P_{s}$.
In Fig. 3(b), unlike the crossover in Fig. 3(a), the $P_{s}$ of the ODM for $N=5\times 10^{3}$ and $10^{4}$ exhibits a plateau within the interval $T_{\delta}/T_{d}\in\left[1.0,1.5\right]$ before approaching $P_{s}=1$ for large $T_{\delta}$. This effect is more apparent in Fig. 3(c), wherein the $P_{s}$ approaches a plateau at $P_{s}\approx 0.75$. To better understand the appearance of such plateaus, we further characterize the switching dynamics of the period-doubled and DTC states using the half winding number $w$, introduced in Ref. [8] to characterize the dynamics of a period-doubled state during the defect. It is defined as [8, 54],
@@
and it quantifies the number of times the $O_{R}(t)$ performs a half-rotation around the origin of the complex plane. To illustrate this for a DTC during a defect protocol, we present in Figs. 4(a)–4(c) the exemplary dynamics of $X=\mathrm{Re}[O_{R}]$ and $Y=\mathrm{Im}[O_{R}]$. Here, we rescale the axes by the maximum value of $|R(t)|$, $R_{m}$. Similar to the results in Ref. [8], and Figs. 1(d), 1(f), and 1(h), during the defect protocol, the trajectory of the $O_{R}$ approaches the origin. The half-winding number depends on whether $O_{R}$ rotates around the origin or not. If $O_{R}$ avoids the origin, then $w=0$, and thus there is no bit-flip. On the other hand, if $O_{R}$ does a single half counter-clockwise rotation around the origin, then $w=+1$, meaning the period-doubled or DTC state switches to its symmetry-broken partner. The same thing occurs when $w=-1$, except that $O_{R}$ now rotates in a clockwise manner. Note that we only get a bit-flip when $w=2n+1$, with $n\in\mathbb{Z}$, since an even-valued $w$ implies a complete rotation around the origin and that the system returns to its original state. Note that our convention for obtaining $\varphi(t)$ using Eq. (6) results to a half-winding number of $w=-1$ in the limit of $T_{\delta}\gg T_{d}$ for both the PO and the ODM.
@@
Figure 4: (a)–(c) Exemplary dynamics of the complex response amplitude of $S^{x}/N$ for (a) $w=1$, (b) $w=0$, and (c) $w=-1$. The corresponding defect durations are (a) $T_{\delta}=0.95T_{d}$, (b) $T_{\delta}=0.5T_{d}$, and (c) $T_{\delta}=1.5$, respectively. (d)–(g) Half-winding number distribution of the resonantly-driven PO for (d) $\tilde{T}=2\times 10^{-4}$, (e) $\tilde{T}=10^{-4}$, (f) $\tilde{T}=10^{-5}$, and (g) $\tilde{T}=10^{-6}$. (h)–(k) Half-winding number distribution of the ODM at resonance for (h) $N=5\times 10^{3}$, (i) $N=10^{4}$, (j) $N=10^{5}$, (k) $N=10^{6}$. The remaining parameters for the PO are $\gamma=0.1\Omega$, and $\delta A=0.4$, while the remaining parameters for the ODM are $\lambda_{0}=0.9\lambda_{c}$, $\kappa=1.0\omega$, and $\delta A=0.1$.
@@
We now show in Figs. 4(d)–4(g) the probability distribution $P_{w}$ of $w$ as a function of $T_{\delta}$ and $\tilde{T}$ for the thermal PO. Note that we have used the same parameters considered in Fig. 3(a). We can see that the $P_{w}$ is localized in a small subset of $w$ when $\tilde{T}\ll 1$. Specifically, for $T_{\delta}<T_{\delta,c}$, $w=0$, and thus no bit flip occurs. Whereas for $T_{\delta}>T_{\delta,c}$, $w=-1$, implying that the period-doubled state switches as a result of the defect protocol. We can also observe that $w$ is discontinuous exactly at $T_{\delta}=T_{\delta,c}$, consistent with the behavior observed in Ref. [8]. However, as we increase $\tilde{T}$, the discontinuity at $T_{\delta,c}$ becomes blurry as $P_{w}$ spreads out such that the PO can either have $w=0$ or $-1$ near $T_{\delta,c}$, i.e., it is now uncertain if the bit-flip operation will work or not. This behavior is consistent with the crossover observed in Fig. 3(a). Note that in the case of $\tilde{T}=2\times 10^{-4}$, as shown in Fig. 4(d), the faint $w=+1$ branch in the $P_{w}$ for $T_{\delta}<T_{\delta,c}$ is due to thermally-activated switching, in which due to the finite coherence time of the oscillations, the thermal noise can become strong enough to flip a period-doubled state [68].
@@
The situation is different in the case of the $P_{w}$ of the ODM, as shown in Figs. 4(h)–4(k). While we still observe qualitatively similar behavior of the distribution as in the PO, such as the localization of $P_{w}$ into one value of $w$ as $N$ increases, the $P_{w}$ of the ODM also reveals a $w=+1$ branch from $T_{\delta}\approx 0.9T_{d}$ to $1.0T_{d}$. In Fig. 4(k), we show that when $N$ is sufficiently large, the $P_{w}$ in these regions of $T_{\delta}$ becomes more concentrated into $w=+1$, explaining the discontinuous transition at $T_{\delta}\approx 0.9T_{d}$ shown in Fig. 3(b). As we decrease $N$, however, the system obtains nonzero probability to either have a half-winding number of $w=\pm 1$ or 00, thus resulting in the plateau observed in Figs. 3(b) and 3(c). Note that while we do not observe any appearance of plateaus in the thermal PO for the parameters considered here, we still expect this behavior to appear when the system has a transition between $w=-1$ and $+1$, and the range of $T_{\delta}$ in which one of the branches exists is small enough for the system to have nonzero probability to get either $w=0$ or ±$\pm 1$.
@@
Our results imply two things about the robustness of the bit-flip operation beyond the weak-dissipation and noiseless limit. First, in the noiseless limit, strong dissipation enhances the robustness of the operation against perturbations in the driving and defect parameters. This leads to a wider range of $A$ and $T_{\delta}$ wherein controlled switching between period-doubled and DTC states can be observed. Second, in the presence of fluctuation for fixed dissipation, not only does the bit-flip operation remains robust for large $T_{\delta}$, but the noise also enhances the switching probability for $T_{\delta}<T_{\delta,c}$. Thus, we demonstrate that the phase ramp as a defect is a robust method for performing bit-flips between the period-doubled states of thermal PO and the DTC states of finite-sized ODM for $\lambda_{0}<\lambda_{c}$.
@@
Given that the DTC states of the ODM for $\lambda_{0}<\lambda_{c}$ can be mapped onto a period-doubled state of a coupled oscillator model [53], we now explore in the next subsection whether we can still switch DTC states using the defect protocol in Eq. (12) when $\lambda_{0}>\lambda_{c}$, in which the mapping of the ODM onto the PO breaks down.